# Inverse problem for semidiscrete stochastic parabolic operators in arbitrary dimensions

### v1

@@ 2. Semidiscrete inverse random source problem
In this section, we establish the conditional stability result for the semidiscrete Cauchy problem, as stated in Theorem 1.3. The section is structured as follows: First, we derive an auxiliary estimate for the solution of a semidiscrete stochastic parabolic equation with non-homogeneous Dirichlet boundary conditions. This estimate plays a crucial role in adapting the Carleman estimate (1.6) to this setting. Second, we present a modified Carleman estimate adapted to the semidiscrete stochastic parabolic operator with nonhomogeneous boundary data, and provide its proof. Finally, we prove Theorem 1.3 by combining the Carleman estimate obtained in the second step with geometric arguments inspired by Yamamoto_2009 .
@@
Let $\xi\in H_{\mathcal{F}}^{1}(0,T;H_{h}^{1/2}(\partial\mathcal{M})\cap H_{h}^{1}(\partial\mathcal{M}))$ with $\xi(0)=0$. Then, for a.e. $\omega\in\Omega$, the solution $u\in L^{2}(0,T;H_{h}^{2}(\mathcal{M}))\cap C(0,T;L_{h}^{2}(\mathcal{M}))$ of the system:
(3.1) $\begin{cases}du-\sum_{i=1}^{n}D_{i}(\gamma_{i}D_{i}u)\,dt=0,&\text{in }\mathcal{M}\times(0,T),\\ u=\xi,&\text{on }\partial\mathcal{M}\times(0,T),\\ u(0)=0,&\text{in }\mathcal{M},\end{cases}$
satisfies the following stability estimate for any constant $C>0$:
(3.2) $\|u\|_{L^{2}_{\mathcal{F}}(0,T;H_{h}^{2}(\mathcal{M}))}+\|u\|_{L^{2}_{\mathcal{F}}(\Omega;C([0,T];L_{h}^{2}(\mathcal{M})))}\leq C\|\xi\|_{H^{1}_{\mathcal{F}}(0,T;H_{h}^{1/2}(\partial\mathcal{M})\cap H_{h}^{1}(\partial\mathcal{M}))}.$
Moreover, the following boundary estimate holds for $t_{r}^{i}(D_{i}u)$ on $\Gamma^{+}$:
(3.3) $\sum_{i=1}^{n}\mathbb{E}\int_{0}^{T}\int_{\Gamma^{+}}t_{r}^{i}(|D_{i}u|^{2})\,dt\leq C\|\xi\|_{H^{1}_{\mathcal{F}}(0,T;H_{h}^{1/2}(\partial\mathcal{M})\cap H_{h}^{1}(\partial\mathcal{M}))}.$
@@
Let $\tilde{u}=u-v$. Since $v(0)=0$ by (3.4) and $\xi(0)=0$, it follows that $\tilde{u}(0)=u(0)-v(0)=0$. Moreover, on $\partial\mathcal{M}\times(0,T)$ we have $\tilde{u}=\xi-\xi=0$. Therefore, $\tilde{u}$ satisfies
@@
(3.5) $\begin{cases}d\tilde{u}-\sum_{i=1}^{n}D_{i}(\gamma_{i}D_{i}\tilde{u})\,dt=\left(\partial_{t}v-\sum_{i=1}^{n}D_{i}(\gamma_{i}D_{i}v)\right)\,dt,&\text{in }\mathcal{M}\times(0,T),\\ \tilde{u}=0,&\text{on }\partial\mathcal{M}\times(0,T),\\ \tilde{u}(0)=0,&\text{in }\mathcal{M}.\end{cases}$
Multiplying the first equation of (3.5) by $\tilde{u}$ and integrating over $\mathcal{M}$, we get
$\int_{\mathcal{M}}\tilde{u}\,d\tilde{u}-\int_{\mathcal{M}}\tilde{u}\sum_{i=1}^{n}D_{i}(\gamma_{i}D_{i}\tilde{u})\,dt=\left(\int_{\mathcal{M}}\tilde{u}\,\partial_{t}v-\int_{\mathcal{M}}\tilde{u}\sum_{i=1}^{n}D_{i}(\gamma_{i}D_{i}v)\right)\,dt.$
Applying integration by parts to the second integral on both sides, we obtain
$\int_{\mathcal{M}}\tilde{u}\,d\tilde{u}+\sum_{i=1}^{n}\int_{\mathcal{M}}\gamma_{i}|D_{i}\tilde{u}|^{2}\,dt=\left(\int_{\mathcal{M}}\tilde{u}\,\partial_{t}v+\sum_{i=1}^{n}\int_{\mathcal{M}}\gamma_{i}D_{i}v\,D_{i}\tilde{u}\right)\,dt,$
where we used $\tilde{u}=0$ on $\partial\mathcal{M}$. Observing that $\tilde{u}\,d\tilde{u}=\frac{1}{2}\,d(\tilde{u}^{2})$ and applying Young’s inequality to the right-hand side of the previous inequality, we get
$\frac{1}{2}\int_{\mathcal{M}}\,d(\tilde{u}^{2})+\frac{1}{2}\sum_{i=1}^{n}\int_{\mathcal{M}}\gamma_{i}|D_{i}\tilde{u}|^{2}\,dt\leq\frac{1}{2}\left(\int_{\mathcal{M}}|\tilde{u}|^{2}+\int_{\mathcal{M}}|\partial_{t}v|^{2}+\sum_{i=1}^{n}\int_{\mathcal{M}}\gamma_{i}|D_{i}v|^{2}\right)\,dt.$
Now, adding $\frac{1}{2}\|\tilde{u}\|_{L^{2}(\mathcal{M})}$ and using (3.4) in the above inequality, we obtain
(3.6) $\frac{1}{2}\,d\left(\|\tilde{u}\|^{2}_{L_{h}^{2}(\mathcal{M})}\right)+\frac{1}{2}\|\tilde{u}\|^{2}_{H_{h}^{1}(\mathcal{M})}\,dt\leq\left(\|\tilde{u}\|^{2}_{L_{h}^{2}(\mathcal{M})}+C\|\xi\|_{H^{1}(0,T;H_{h}^{1/2}(\partial\mathcal{M}))}\right)\,dt.$
Letting $E(t)=\|\tilde{u}\|^{2}_{L^{2}_{h}(\mathcal{M})}$ and $\mathcal{C}_{1}=C\|\xi\|_{H^{1}(0,T;H_{h}^{1/2}(\partial\mathcal{M}))}$ in (3.6), we have
$\frac{1}{2}\,dE(t)\leq(E(t)+\mathcal{C}_{1})\,dt.$
Multiplying by $e^{-2t}$ and using Itô’s lemma, we obtain
$\begin{split}e^{-2t}\,dE(t)-e^{-2t}E(t)\,dt\leq&e^{-2t}\mathcal{C}_{1}\,dt,\\ d\left(e^{-2t}E(t)\right)\leq&e^{-2t}\mathcal{C}_{1}\,dt.\end{split}$
Since $E(0)=0$, integrating over $(0,t)$ for $t\in[0,T]$, we have
(3.7) $\|\tilde{u}\|^{2}_{L_{h}^{2}(\mathcal{M})}=E(t)\leq\int_{0}^{t}e^{2(t-s)}\mathcal{C}_{1}\,ds=\frac{-1+e^{2t}}{2}\mathcal{C}_{1},$
for all $t\in[0,T]$. Recalling that $u=\tilde{u}+v$, and using (3.4) and (3.7), we get
$\|u\|^{2}_{L^{2}_{h}(\mathcal{M})}\leq\|\tilde{u}\|^{2}_{L_{h}^{2}(\mathcal{M})}+\|v\|^{2}_{L_{h}^{2}(\mathcal{M})}\leq\frac{e^{2t}-1}{2}\mathcal{C}_{1}+C\|\xi\|_{H^{1}(0,T;H_{h}^{1/2}(\partial\mathcal{M}))}.$
@@
Therefore, taking the supremum over $[0,T]$ in the above inequality, we conclude that
$\|u\|_{C(0,T;L_{h}^{2}(\mathcal{M}))}\leq C\|\xi\|_{H^{1}(0,T;H_{h}^{1/2}(\partial\mathcal{M}))}.$
Furthermore, using (3.6), (3.7) and integrating over $(0,T)$, we have
$\|u\|_{L_{h}^{2}(0,T;H^{1}(\mathcal{M}))}\leq C\|\xi\|_{H_{h}^{1}(0,T;H^{1/2}(\partial\mathcal{M}))}.$
@@
Combining the last two inequalities and taking the expectation, we derive the first estimate (3.2).
In turn, we note that multiplying by $x_{j}A_{j}D_{j}u$, where $x_{j}:=e_{j}\cdot x$, the main equation from (3.1) we obtain
(3.8) $du\,x_{j}A_{j}D_{j}u=\sum_{i=1}^{n}D_{i}(\gamma_{i}D_{i}u)x_{j}D_{j}A_{j}udt.$
First, let us focus on the right-hand side above. Thanks to (1.13) we write
(3.9) $\sum_{i=1}^{n}D_{i}(\gamma_{i}D_{i}u)x_{j}A_{j}D_{j}udt=$ $x_{j}D_{j}\gamma_{j}|A_{j}D_{j}u|^{2}+x_{j}A_{j}\gamma_{j}D_{j}^{2}u\,D_{j}A_{j}u$
$+\sum_{i\neq j}x_{j}D_{i}(\gamma_{i}D_{i}u)\,D_{j}A_{j}u.$
@@
Then, writing $D_{j}^{2}uD_{j}A_{j}u=\frac{1}{2}D_{j}(|D_{j}u|^{2})$, integrating the results over $G\times(0,T)$, and applying an integration by parts it follows
$\frac{1}{2}\int_{\mathcal{M}}x_{j}A_{j}\gamma_{j}D_{j}(|D_{j}u|^{2})=$ $-\frac{1}{2}\int_{\mathcal{M}^{\ast}_{j}}D_{j}(A_{j}\gamma_{j}x_{j})|D_{j}u|^{2}$
$+\frac{1}{2}\int_{\partial_{j}\mathcal{M}}x_{j}A_{j}\gamma_{j}t_{r}^{j}(|D_{j}u|^{2})\nu_{j}.$
@@
Thus, denoting $\Gamma_{j}:=\{x\in\partial_{j}\mathcal{M}\mid x_{j}\nu_{j}=1\}\subset\mathcal{M}_{j}$ it follows that the above equals
(3.10) $\frac{1}{2}\int_{\mathcal{M}}x_{j}A_{j}\gamma_{j}D_{j}(|D_{j}u|^{2})=-\frac{1}{2}\int_{\mathcal{M}^{\ast}_{j}}D_{j}(A_{j}\gamma_{j}x_{j})|D_{j}u|^{2}+\frac{1}{2}\int_{\Gamma_{j}}A_{j}\gamma_{j}t_{r}^{j}(|D_{j}u|^{2}).$
@@
(3.11) $\int_{G}\,x_{j}A_{j}D_{j}u\,du=$ $\int_{G}x_{j}D_{j}\gamma_{j}\,|A_{j}D_{j}u|^{2}-\frac{1}{2}\int_{\mathcal{M}_{j}^{\ast}\times(0,T)}D_{j}(A_{j}(\gamma_{j})x_{j})|D_{j}u|^{2}$
$+\frac{1}{2}\int_{\Gamma_{j}\times(0,T)}A_{j}\gamma_{j}t_{r}^{j}(|D_{j}u|^{2})+\sum_{i\neq j}\int_{G}x_{j}D_{i}(\gamma_{i}\,D_{i}u)\,D_{j}A_{j}u.$
Finally, using (3.8) in the integral of the left-hand side above, that $\gamma_{i}$ is bounded for each $j\in\llbracket 1,n\rrbracket$ and the Young inequality we conclude that there exists a constant $C$ independent of $h$ such that
(3.12) $\int_{\Gamma_{j}\times(0,T)}t_{r}^{j}(|D_{j}u|^{2})dt\leq C\left\|u\right\|^{2}_{L^{2}(0,T;H_{h}^{2}(\mathcal{M}))}.$
@@
(3.13) $\sum_{j=1}^{n}\int_{\Gamma_{j}\times(0,T)}t_{r}^{j}(|D_{j}u|^{2})dt\leq C\left\|u\right\|^{2}_{L^{2}(0,T;H_{h}^{2}(\mathcal{M}))}.$
@@
3.1. Semidiscrete Carleman estimate with nonhomogeneous Dirichlet condition.
In the Carleman estimate (1.10), the function $\psi$ is chosen as $|x-x^{\ast}|^{2}$. More generally, we may consider the following
(3.14) $\psi(x)=\Psi(x),$
where $|\nabla\Psi|>0$ in $G$, so that the assumptions required to obtain the Carleman estimate (1.10) remain valid. This generalized setting will later allow us to derive a Carleman estimate for a stochastic parabolic operator with non-homogeneous Dirichlet boundary conditions.
Based on this choice of $\psi$, we now state the corresponding Carleman estimate, which extends (1.10) to the present setting. For completeness, we also provide its proof in the following.

### v2

@@ 2. Semi-discrete inverse random source problem
In this section, we establish the conditional stability result for the semi-discrete Cauchy problem, as stated in Theorem 1.3. The section is structured as follows: First, we introduce a new weight function in the semi-discrete context to derive a new semi-discrete Carleman estimate with boundary data, using a similar procedure for obtaining (1.5). Second, we need an auxiliary estimate for the solution of a semi-discrete stochastic parabolic equation with non-homogeneous Dirichlet boundary conditions. This estimate plays a crucial role in the proof a new Carleman estimate (1.11) to this setting. Second, we present a modified Carleman estimate adapted to the semi-discrete stochastic parabolic operator with nonhomogeneous boundary data, and provide its proof. Finally, we prove Theorem 1.3 by combining the Carleman estimate obtained in the second step with geometric arguments inspired by Yamamoto (2009).
We begin by deriving the new weight function required for the semi-discrete Carleman estimate with boundary data. Following the strategy used in Boyer et al. (2010b) to obtain Assumption 2.1, and adapting the continuous construction of (Èmanuilov, 1995, Lemma 1.2) to the semi-discrete framework, we obtain the existence of such a function as follows:
Assumption 3.1.
[…]
$\begin{array}[]{ccc}d_{0}(x)>0\quad\text{in}\,\,\tilde{G},&\quad|\nabla d_{0}(x)|>c\quad\text{in}\,\,x\in\overline{\tilde{G}},&\sum_{i=1}^{n}\partial_{i}d_{0}(x)\nu_{i}(x)\leq 0,\quad x\in V_{\partial_{i}G\setminus\Gamma},\end{array}$
where $V_{\partial_{i}G\setminus\Gamma}$ is a sufficiently small neighborhood of $\partial_{i}G\setminus\Gamma$ in $\tilde{G}\setminus\tilde{\Gamma}$, in which the outward normal unit $\nu_{i}$ to $G$ extends from $\partial_{i}G$.
Example 3.2.
Let $x^{*}\notin\overline{G}$. We define the set $\mathcal{M}$ and
$\Gamma:=\left\{x\in\partial\mathcal{M}:x\in\partial_{i}\mathcal{M}\ \text{and}\ (x-x^{*})\cdot e_{i}\,\nu_{i}>0\right\}.$
For these sets, a function that satisfies the assumption 3.1 is $d_{0}(x)=|x-x^{*}|^{2}.$
In this construction, the size of the observation set $\Gamma$ depends on the choice of the reference point $x^{*}$. Indeed, in dimension $2$, let us consider a point $x^{*}=(a,b)$ with $a<0$ and $0<b<1$. In this case, if we consider $L:=(0,1)\cap h\mathbb{Z}$, then $\Gamma=\{(x,0)|\,x\in L\}\cup\{(1,x)|\,x\in L\}\cup\{(x,1)|\,x\in L\}.$ If instead $b>1$, then $\Gamma=\{(x,0)|\,x\in L\}\cup\{(1,x)|\,x\in L\},$ while for $b<0$ we have $\Gamma=\{(1,x)|\,x\in L\}\cup\{(x,1)|\,x\in L\}.$ Therefore, the set $\Gamma$ contains more faces of the square when $0<b<1$ than when $b>1$ or $b<0$. In general, if $x^{*}$ is chosen so that there exists some $x\in\partial\mathcal{M}$ for which the vector $x-x^{*}$ is orthogonal to a face of $\partial\mathcal{M}$, then the corresponding portion of $\Gamma$ becomes larger. In the opposite case, the set $\Gamma$ becomes smaller. Thus, a suitable choice of $x^{*}$ can be used to minimize the required observation region.
The following result may be proved in much the same way as Theorem 1.5.
Theorem 3.3.
Let $d_{0}$ satisfy assumptions 3.1 and $\varphi=e^{\lambda d_{0}}$. For $\lambda_{1}\geq 1$ sufficiently large, there exist $C$, $\tau_{0}\geq 1$, $h_{0}>0$, $\varepsilon_{0}>0$, depending on $G$, $\beta$, $c_{0}$, $T$, and $\lambda$, such that
(3.1) $\begin{split}&J(w)+\mathbb{E}\int_{Q}s^{3}\lambda^{4}_{1}\varphi^{3}\,e^{2s\varphi}|w|^{2}\,dt+\sum_{i,j=1}^{n}\mathbb{E}\int_{Q_{ij}^{\ast}}s^{-1}e^{2s\varphi}|D_{ij}^{2}w|^{2}\,dt+\mathbb{E}\int_{Q}s\lambda_{1}^{2}\varphi\,e^{2s\varphi}|g|^{2}dt\\ &\leq\,\mathcal{C}_{\tau_{0},\varepsilon_{0}}\left(\mathbb{E}\int_{Q}\,e^{2s\varphi}|f|^{2}\,dt+\sum_{i=1}^{n}\mathbb{E}\int_{Q_{i}^{\ast}}\,e^{2s\varphi}|D_{i}g|^{2}dt+\sum_{i=1}^{n}\mathbb{E}\int_{0}^{T}\int_{\Gamma}st_{r}^{i}(e^{2s\varphi}|D_{i}w|^{2})\,dt\right.\\ &\left.+\left.\mathbb{E}\int_{\mathcal{M}}s^{2}\,e^{2s\varphi}|w|^{2}\right|_{t=T}\right).\end{split}$
for all $\tau\geq\tau_{0}$, $0<h\leq h_{0}$, $s(t)h\leq\epsilon_{0}$, where $f\in L^{2}_{\mathcal{F}}(0,T;L^{2}(\mathcal{M}))$, $g\in L^{2}_{\mathcal{F}}(0,T;H^{1}(\mathcal{M}))$, and $w$ satisfy $dw-\sum_{i=1}^{n}D_{i}(\gamma_{i}D_{i}w)\,dt=fdt+gdB(t)$ with $w=0$ on $\partial\mathcal{M}$ and $w|_{t=0}=0$ in $\mathcal{M}$.
Proof.
We only outline the main ideas of the proof. If we replace $\psi(x)$ by $d_{0}(x)$ given in assumption (3.1) in Appendix A, all computations remain unchanged, except for the estimate in (A.11), due to the different properties of the new weight function. Thus, in (A.11), the terms supported in $G_{0}\cap\mathcal{M}$ are replaced by boundary trace terms of the derivatives on a portion of the boundary. Indeed, for all $\tau\geq\tau_{1}>1$, $0<h<h_{0}$, and $s(t)h\leq\varepsilon_{1}$, we obtain
(3.2) $\mathbb{E}\int_{Q}$ $s^{3}\lambda^{4}_{1}\varphi^{3}|z|^{2}\,dt+\mathbb{E}\int_{Q}(s^{2}\lambda_{1}^{2}\varphi^{2}+s\lambda_{1}^{2}\varphi)|dz|^{2}+\mathbb{E}\int_{Q}|Cz|^{2}\,dt$
$+\sum_{i=1}^{n}\left(\mathbb{E}\int_{Q_{i}^{\ast}}s\lambda^{2}_{1}\varphi\,|D_{i}z|^{2}\,dt+\mathbb{E}\int_{Q}s\lambda_{1}^{2}\varphi|A_{i}D_{i}z|^{2}\,dt-\mathbb{E}\int_{Q_{i}^{\ast}}|D_{i}(dz)|^{2}\right)$
$\leq\mathcal{C}_{\tau_{0},\varepsilon_{0}}\left(\mathbb{E}\int_{Q}|rf|^{2}\,dt+\sum_{i=1}^{n}\mathbb{E}\int_{0}^{T}\int_{\Gamma}st_{r}^{i}(|D_{i}z|^{2})\,dt+\left.\mathbb{E}\int_{\mathcal{M}}s^{2}|z|^{2}\right|_{t=T}\right),$
where $z=e^{-s\varphi}w$, and $w$ satisfies $dw-\sum_{i=1}^{n}D_{i}(\gamma_{i}D_{i}w)\,dt=f\,dt+g\,dB(t).$
We now return to the original variable $w$. The argument used in Section A.3 applies with minor modifications. The only change is in the right-hand side of (3.2), which must be adapted accordingly; the details are given in Appendix C. This yields the Carleman estimate with boundary data (3.1). ∎
From (3.1), it may be conclude a result of Lipschitz stability for the semi-discrete inverse random source problem with boundary data that generalizes the result found in Wu et al. (2024) to arbitrary dimensions as
Theorem 3.4.
Let $a_{1i}\in L^{\infty}_{\mathcal{F}}(0,T;L_{h}^{\infty}(\mathcal{M}))$ for $i=1,\ldots,n$, and $a_{2}\in L_{\mathcal{F}}^{\infty}(0,T;L_{h}^{n^{\ast}}(\mathcal{M}))$ where $n^{\ast}$ satisfies (2.4). Assume that $g_{k}\in L^{2}_{\mathcal{F}}(0,T;H_{h}^{1}(\mathcal{M}))$ and satisfies the condition (1.6) for $k=1,2$. Then, there exists a constant $\mathcal{C}>0$, independent of $h$, such that
(3.3) $\|g_{1}-g_{2}\|_{L^{2}_{\mathcal{F}}(0,T;L_{h}^{2}(\mathcal{M}))}\leq\mathcal{C}\|\Lambda_{3}(g_{1})-\Lambda_{3}(g_{2})\|_{(L^{2}_{\mathcal{F}}(0,T;L^{2}_{h}(\Gamma)))^{n}\times L^{2}_{\mathcal{F}_{T}}(\Omega,L^{2}_{h}(\mathcal{M}))}.$
where $\Lambda_{3}$ is defined as
$\Lambda_{3}(g)=(\left.\partial_{\nu}w\right|_{(0,T)\times\Gamma},\left.w\right|_{t=T})$
for $w$ solution of (1.4).
We continue in this fashion obtaining the following result about the estimate for the solution of a semi-discrete homogeneous stochastic parabolic equation with non-homogeneous Dirichlet boundary conditions.
@@
Let $\xi\in H_{\mathcal{F}}^{1}(0,T;H_{h}^{1/2}(\partial\mathcal{M})\cap H_{h}^{1}(\partial\mathcal{M}))$ with $\left.\xi\right|_{t=0}=0$. Then, for a.e. $\omega\in\Omega$, the solution $u\in L^{2}(0,T;H_{h}^{2}(\mathcal{M}))\cap C(0,T;L_{h}^{2}(\mathcal{M}))$ of the system:
(3.4) $\begin{cases}du-\sum_{i=1}^{n}D_{i}(\gamma_{i}D_{i}u)\,dt=0,&\text{in }Q,\\ u=\xi,\quad\text{on }\partial Q,\quad\left.u\right|_{t=0}=0,\quad\text{in }\mathcal{M},\end{cases}$
satisfies the following stability estimate for any constant $\mathcal{C}>0$:
(3.5) $\|u\|_{L^{2}_{\mathcal{F}}(0,T;H_{h}^{2}(\mathcal{M}))}+\|u\|_{L^{2}_{\mathcal{F}}(\Omega;C([0,T];L_{h}^{2}(\mathcal{M})))}\leq\mathcal{C}\|\xi\|_{H^{1}_{\mathcal{F}}(0,T;H_{h}^{1/2}(\partial\mathcal{M})\cap H_{h}^{1}(\partial\mathcal{M}))}.$
Moreover, the following boundary estimate holds for $t_{r}^{i}(D_{i}u)$ on $\partial\mathcal{M}$:
(3.6) $\sum_{i=1}^{n}\mathbb{E}\int_{\partial Q}t_{r}^{i}(|D_{i}u|^{2})\,dt\leq\mathcal{C}\|\xi\|_{L^{2}_{\mathcal{F}}(0,T;H_{h}^{1/2}(\partial\mathcal{M})\cap H_{h}^{1}(\partial\mathcal{M}))}.$
@@
Let $\tilde{u}=u-v$. Since $v|_{t=0}=0$ by (3.7) and $\xi|_{t=0}=0$, it follows that $\tilde{u}|_{t=0}=u|_{t=0}-v|_{t=0}=0$. Moreover, on $\partial\mathcal{M}\times(0,T)$ we have $\tilde{u}=\xi-\xi=0$. Therefore, $\tilde{u}$ satisfies
@@
(3.8) $\begin{cases}d\tilde{u}-\sum_{i=1}^{n}D_{i}(\gamma_{i}D_{i}\tilde{u})\,dt=\left(\partial_{t}v-\sum_{i=1}^{n}D_{i}(\gamma_{i}D_{i}v)\right)\,dt,&\text{in }\mathcal{M}\times(0,T),\\ \tilde{u}=0,\quad\quad\text{on }\partial\mathcal{M}\times(0,T),\quad\quad\left.\tilde{u}\right|_{t=0}=0,\quad\text{in }\mathcal{M}.\end{cases}$
The task is now to find an energy estimate for $u$. Let us apply the Itô’s formula to $\frac{1}{2}\tilde{u}^{2}$ yields $\frac{1}{2}d(\tilde{u}^{2})=\tilde{u}d\tilde{u}+\frac{1}{2}|d\tilde{u}|^{2}$. As $\tilde{u}$ satisfies (3.8), we have
$\frac{1}{2}d(\tilde{u}^{2})=\sum_{i=1}^{n}\tilde{u}D_{i}(\gamma_{i}D_{i}\tilde{u})\,dt+\tilde{u}\partial_{t}v\,dt-\sum_{i=1}^{n}\tilde{u}D_{i}(\gamma_{i}D_{i}v)\,dt$
By integration over $\mathcal{M}$ and integration by parts to the first and third terms, we get
$\frac{1}{2}\int_{\mathcal{M}}d(\tilde{u}^{2})=-\sum_{i=1}^{n}\int_{\mathcal{M}_{i}^{\ast}}\gamma_{i}|D_{i}(\tilde{u})|^{2}\,dt+\int_{\mathcal{M}}\tilde{u}\partial_{t}v\,dt+\sum_{i=1}^{n}\int_{\mathcal{M}_{i}^{*}}\gamma_{i}D_{i}(\tilde{u})\;D_{i}(v)\,dt,$
which is due to the fact that $\tilde{u}=0$ on $\partial\mathcal{M}$. From Young’s inequality on the second and third terms in the above equation, it follows that
$\frac{1}{2}\int_{\mathcal{M}}d(\tilde{u}^{2})\leq-\frac{1}{2}\sum_{i=1}^{n}\int_{\mathcal{M}_{i}^{*}}\gamma_{i}|D_{i}\tilde{u}|^{2}\,dt+\frac{1}{2}\int_{\mathcal{M}}|\tilde{u}|^{2}\,dt+\frac{1}{2}\int_{\mathcal{M}}|\partial_{t}v|^{2}\,dt+\frac{1}{2}\sum_{i=1}^{n}\int_{\mathcal{M}_{i}^{*}}\gamma_{i}|D_{i}v|^{2}\,dt$
Now, integrating on $[0,t]$ for $t\in[0,T]$, using (3.7) of the above inequality and from condition that $\left.\tilde{u}\right|_{t=0}=0$, we can assert that
(3.9) $\frac{1}{2}\,\int_{\mathcal{M}}|\tilde{u}|^{2}+\frac{1}{2}\sum_{i=1}^{n}\int_{Q_{i}^{*}}\gamma_{i}|D_{i}\tilde{u}|^{2}\,dt\leq\mathcal{C}\int_{0}^{T}\left(\|\xi_{t}\|^{2}_{H_{h}^{1/2}(\partial\mathcal{M})}+\|\xi\|^{2}_{H_{h}^{1/2}(\partial\mathcal{M})}\right)\,dt.$
Recalling $u=\tilde{u}+v$ and using the above inequality, we deduce that
$\frac{1}{2}\int_{\mathcal{M}}|u|^{2}+\frac{1}{2}\sum_{i=1}^{n}\int_{Q_{i}^{*}}\gamma_{i}|D_{i}u|^{2}\,dt\leq\mathcal{C}\int_{0}^{T}\left(\|\xi_{t}\|^{2}_{H_{h}^{1/2}(\partial\mathcal{M})}+\|\xi\|^{2}_{H_{h}^{1/2}(\partial\mathcal{M})}\right)\,dt+\frac{1}{2}\int_{\mathcal{M}}|v|^{2}.$
To estimate the term $\int_{\mathcal{M}}|v|^{2}$, we note that applying Young’s inequality and (3.7) enables us to write
$\frac{1}{2}\int_{\mathcal{M}}\partial_{t}|v|^{2}\leq\frac{1}{2}\int_{\mathcal{M}}v\,\partial_{t}v\leq\frac{1}{2}\int_{\mathcal{M}}|v|^{2}+|v_{t}|^{2}\leq\frac{1}{2}\left(\|\xi\|^{2}_{H^{1/2}_{h}(\partial\mathcal{M})}+\|\xi_{t}\|_{H^{1/2}_{h}(\partial\mathcal{M})}\right)$
Thus,
$\int_{\mathcal{M}}|v|^{2}\leq\int_{0}^{T}\|\xi_{t}\|^{2}_{H_{h}^{1/2}(\partial\mathcal{M})}+\|\xi\|^{2}_{H_{h}^{1/2}(\partial\mathcal{M})}\,dt$
where we used the integration on $[0,T]$ and that $\left.v\right|_{t=0}=0$. Combining these inequalities, we conclude that
(3.10) $\frac{1}{2}\int_{\mathcal{M}}|u|^{2}+\frac{1}{2}\sum_{i=1}^{n}\int_{Q_{i}^{*}}\gamma_{i}|D_{i}u|^{2}\,dt\leq\mathcal{C}\int_{0}^{T}\left(\|\xi_{t}\|^{2}_{H_{h}^{1/2}(\partial\mathcal{M})}+\|\xi\|^{2}_{H_{h}^{1/2}(\partial\mathcal{M})}\right)\,dt.$
@@
Therefore, taking the supremum over $[0,T]$ in the above inequality only considered the first terms, we conclude that
$\|u\|_{C(0,T;L_{h}^{2}(\mathcal{M}))}\leq\mathcal{C}\int_{0}^{T}\|\xi_{t}\|^{2}_{H^{1/2}_{h}(\partial\mathcal{M})}+\|\xi\|^{2}_{H_{h}^{1/2}(\partial\mathcal{M})}dt.$
Furthermore, using (3.10), the above inequality and integrating over $(0,T)$, we have
$\|u\|_{L_{h}^{2}(0,T;H^{1}_{h}(\mathcal{M}))}\leq\mathcal{C}\int_{0}^{T}\|\xi_{t}\|^{2}_{H^{1/2}_{h}(\partial\mathcal{M})}+\|\xi\|^{2}_{H_{h}^{1/2}(\partial\mathcal{M})}dt.$
Similar arguments apply to the case $D_{i}u$ for each $i=1,\ldots,n$, yields
$\|u\|_{L_{h}^{2}(0,T;H_{h}^{2}(\mathcal{M}))}\leq\mathcal{C}\int_{0}^{T}\|\xi_{t}\|^{2}_{H^{1/2}_{h}(\partial\mathcal{M})\cap H^{1}(\partial\mathcal{M})}+\|\xi\|^{2}_{H_{h}^{1/2}(\partial\mathcal{M})\cap H^{1}_{h}(\partial\mathcal{M})}dt.$
@@
Combining these inequalities and taking the expectation, we derive the first estimate (3.5).
Our next goal is to determine the estimate for $\sum_{i=1}^{n}\mathbb{E}\int_{\partial Q}t_{r}^{i}(|D_{i}u|^{2})\,dt$. In order to obtain this inequality, it is convenient to consider the following
(3.11) $2x_{j}D_{j}^{2}(u)A_{j}D_{j}u=x_{j}D_{j}(|D_{j}u|^{2})$
where $x_{j}=x\cdot e_{j}$.
@@
Integrating the results over $\mathcal{M}\times(0,T)$ and applying an integration by parts it follows
$\frac{1}{2}\int_{\mathcal{M}}x_{j}D_{j}(|D_{j}u|^{2})=$ $-\frac{1}{2}\int_{\mathcal{M}^{\ast}_{j}}D_{j}(x_{j})|D_{j}u|^{2}+\frac{1}{2}\int_{\partial_{j}\mathcal{M}}x_{j}t_{r}^{j}(|D_{j}u|^{2})\nu_{j}.$
@@
Thus, denoting $\Gamma_{j}:=\{x\in\partial_{j}\mathcal{M}\mid x_{j}\nu_{j}=1\}\subset\partial_{j}\mathcal{M}$ it follows that the above equals
(3.12) $\frac{1}{2}\int_{\mathcal{M}}x_{j}D_{j}(|D_{j}u|^{2})=-\frac{1}{2}\int_{\mathcal{M}_{j}^{\ast}}|D_{j}u|^{2}+\frac{1}{2}\int_{\Gamma_{j}}t_{r}^{j}(|D_{j}u|^{2}).$
@@
(3.13) $\int_{Q}\,x_{j}D_{j}^{2}(u)A_{j}D_{j}(u)\,\,dt=-\frac{1}{2}\int_{\mathcal{M}_{j}^{\ast}}|D_{j}u|^{2}+$ $\frac{1}{2}\int_{\Gamma_{j}\times(0,T)}t_{r}^{j}(|D_{j}u|^{2})\,dt$
Finally, using the Young inequality and taking expectation we conclude that there exists a constant $\mathcal{C}$ independent of $h$ such that
(3.14) $\mathbb{E}\int_{\Gamma_{j}\times(0,T)}t_{r}^{j}(|D_{j}u|^{2})dt\leq\mathcal{C}\left\|u\right\|^{2}_{L^{2}_{\mathcal{F}}(0,T;H_{h}^{2}(\mathcal{M}))}.$
@@
(3.15) $\sum_{i=1}^{n}\mathbb{E}\int_{\partial Q}t_{r}^{i}(|D_{i}u|^{2})\,dt\leq\mathcal{C}\left\|u\right\|^{2}_{L_{\mathcal{F}}^{2}(0,T;H_{h}^{2}(\mathcal{M}))}.$
@@
3.1. Semi-discrete Carleman estimate with nonhomogeneous Dirichlet condition.
Based on this choice of $\psi=d_{0}$, we now state the corresponding Carleman estimate, which extends (3.1) to the present setting. For completeness, we also provide its proof in the following.